OpenStudy (anonymous):

2/5(9-2b)=-14 ---how to solve and find what's "b" step by step?

OpenStudy (anonymous):

is it ??/ 2 -- (9 - 2b) = -14 5

OpenStudy (anonymous):

b=-81.5

OpenStudy (anonymous):

oh--- 2/5 is the fraction 2 over 5

OpenStudy (anonymous):

and no, the answer isn't 81.5, it didn't worked out :P .......

OpenStudy (anonymous):

If yes, then multiply both sides by 5/2 5 2 5 --- * --- (9 - 2b) = --- * -14 2 5 2 9 - 2b = -35 9 + 35 = 2b 44 = 2b 44/2 = b 22 = b